OpenStudy (anonymous):

Write an equation for the horizontal line that contains point E(6, –6). a. x=6 b. x=-6 c. y=-6 d. y=6

OpenStudy (anonymous):

when x = 6, y = -6 when y = -6, x = 6 therefore, y = x -12

OpenStudy (anonymous):

lets try when x = 6: y = (6) -12 y= -6 as in point E (6, -6) when y = -6: -6 = x -12 -x = 6 -12 -x = -12 x = 6 as in point E